### v1

@@ Abstract
In this paper, we propose an adaptive keyframe selection method for improved 3D scene reconstruction in dynamic environments. The proposed method integrates two complementary modules: an error-based selection module utilizing photometric and structural similarity (SSIM) errors, and a momentum-based update module that dynamically adjusts keyframe selection thresholds according to scene motion dynamics. By dynamically curating the most informative frames, our approach addresses a key data bottleneck in real-time perception. This allows for the creation of high-quality 3D world representations from a compressed data stream, a critical step towards scalable robot learning and deployment in complex, dynamic environments. Experimental results demonstrate significant improvements over traditional static keyframe selection strategies, such as fixed temporal intervals or uniform frame skipping. These findings highlight a meaningful advancement toward adaptive perception systems that can dynamically respond to complex and evolving visual scenes. We evaluate our proposed adaptive keyframe selection module on two recent state-of-the-art 3D reconstruction networks, Spann3r and CUT3R, and observe consistent improvements in reconstruction quality across both frameworks. Furthermore, an extensive ablation study confirms the effectiveness of each individual component in our method, underlining their contribution to the overall performance gains.
@@ 3.2.1 Photometric Error
where $p$ represents a pixel coordinate in the valid mask $\mathcal{M}$, and $I(p)$ is the pixel intensity. We specifically choose the L1 norm over the L2 norm (Mean Squared Error) for its increased robustness to outliers, such as sensor noise or small occlusions, which might otherwise disproportionately influence the error score.
@@ 3.2.3 Combined Error Score
1:function ComputeHybridError($I_{t},I_{k},\text{Pose}_{t},\text{Pose}_{k},\text{Intrinsics}$)
2:  $\hat{I}_{k},\mathcal{M}\leftarrow\texttt{WarpFrame}(I_{k},\text{Pose}_{k},\text{Pose}_{t},\text{Intrinsics})$
@@ 3.4 Hyperparameter Analysis
The core logic of our error computation is summarized in Algorithm 1. The behavior of our framework is governed by a set of key hyperparameters. In our experiments, these were set as follows based on empirical evaluation on a held-out validation dataset:
  • Error Weights ($\alpha,\beta$): Set to $\alpha=0.7$ and $\beta=0.3$ to prioritize the capture of fine-grained detail.
  • Window Size ($W$): A window size of $W=5$ frames was used. This provides a balance between statistical stability and reactivity to sudden changes in scene dynamics.
@@
  • Base Threshold ($\theta_{0}$): A small base threshold was set to ensure a minimum level of selectivity even in completely static scenes.
@@ 4.1 Datasets and Evaluation Metrics
We evaluate reconstruction quality using a suite of standard metrics as in prior works [4, 5, 7, 6]: Accuracy (Acc↓) and Completion (Comp↓) measure the distance between the predicted and ground truth point clouds, Normal Consistency (NC↑) assesses surface smoothness, and Chamfer Distance (CD↓) captures overall shape similarity. To specifically quantify selection efficiency, we introduce the Keyframe Compression Ratio (KFCR↑). This metric represents the percentage of frames discarded from the original sequence, where a higher value indicates greater efficiency and compression.
@@ 4.2.2 Performance on Dynamic Datasets
On the challenging real-world Bonn dataset, our method achieves superior results across all reconstruction metrics for both Spann3r and CUT3R, while also using fewer keyframes (higher KFCR). This demonstrates the core advantage of our approach: by intelligently selecting frames that capture meaningful changes, it not only improves efficiency but also leads to a more accurate and complete final reconstruction. The ability to filter redundant information allows the reconstruction network to focus its capacity on modeling the most critical aspects of the evolving scene.
@@ 4.3 Qualitative Analysis
Figure 2: Qualitative Results on the dynamic BONN dataset. We compare our method with concurrent works, CUT3R. Our method achieves the best qualitative results in complex and cluttered environments.
Beyond quantitative metrics, we provide a qualitative analysis to inspect the differences in reconstruction quality visually. As illustrated in Figure 2, which shows results on challenging sequences from the BONN dataset, our method produces visibly more complete and detailed 3D models. When integrated with networks such as CUT3R [6], our adaptive selection helps capture fine details on moving objects and reduces artifacts in the reconstructed geometry. The impact of the module is more significant when evaluated on outdoor scenes, particularly in dynamic environments. This visual evidence supports the quantitative findings and highlights the practical benefits of our approach to real-time robotic perception in dynamic environments.
@@ 4.4.1 Ablation Study
In contrast, our full model integrating both photometric and SSIM errors achieves the best balance of reconstruction quality and efficiency. It outperforms all baselines in accuracy, Chamfer distance, and normal consistency, while maintaining a competitive keyframe compression ratio. This confirms that the complementary integration of photometric sensitivity and structural robustness yields superior keyframe selection performance compared to prior strategies.
@@ 4.4.2 Scene-wise Efficiency and Quality
5 CONCLUSIONS AND FUTURE WORK
@@
For future work, our work on efficient world understanding serves as a foundational step for several exciting research directions. First, while our hybrid error metric is robust, its reliance on appearance makes it sensitive to challenging illumination conditions [25], which we want to address in the future. Second, we plan to leverage the high-fidelity and compact 3D representations created by our method to address the long-tail data problem in robotics. The curated keyframes and resulting 3D models can serve as grounded "digital twins" of real-world events.

### v3

@@ Abstract
In this paper, we propose an adaptive keyframe selection method for improved 3D scene reconstruction in dynamic environments using RGB-D sensors. The proposed method integrates two complementary modules: an error-based selection module utilizing photometric and structural similarity (SSIM) errors derived from depth-based warping, and a momentum-based update module that dynamically adjusts keyframe selection thresholds according to scene motion dynamics. By dynamically curating the most informative frames, our approach addresses a key data bottleneck in real-time perception. This allows for the creation of high-quality 3D world representations from a compressed data stream, a critical step towards scalable robot learning and deployment in complex, dynamic environments. Experimental results demonstrate significant improvements over traditional static keyframe selection strategies, such as fixed temporal intervals or uniform frame skipping. These findings highlight a meaningful advancement toward adaptive perception systems that can dynamically respond to complex and evolving visual scenes. We evaluate our proposed adaptive keyframe selection module on two recent state-of-the-art 3D reconstruction networks, Spann3r and CUT3R, and observe consistent improvements in reconstruction quality across both frameworks. Furthermore, an extensive ablation study confirms the effectiveness of each individual component in our method, underlining their contribution to the overall performance gains.
@@ 1 Introduction
While classical SLAM systems like ORB-SLAM and DSO utilize dynamic keyframe insertion strategies based on tracking quality or parallax, these are primarily optimized for localization stability. In contrast, dense reconstruction often defaults to static temporal intervals or uniform skipping to manage data rates. We distinguish our work by focusing on reconstruction quality rather than tracking. Keyframe selection can be broadly categorized into static (fixed interval) [10, 4, 5, 7, 6], heuristic (motion-based) [11], geometric (parallax/overlap) [14], and appearance-based (photometric) methods [ours]. Our approach bridges these by combining geometric warping with appearance-based metrics to optimize the density of the 3D representation.
@@ 3.2.1 Photometric Error
where $p$ represents a pixel coordinate in the valid mask $\mathcal{M}$, and $I(p)$ is the pixel intensity. We specifically choose the L1 norm over the L2 norm (Mean Squared Error) for its increased robustness to outliers, such as sensor noise or small occlusions, which might otherwise disproportionately influence the error score. The valid mask $\mathcal{M}$ identifies pixels with successful depth-based reprojection between frames. We observe that as the overlap decreases, common in highly dynamic sequences, the mask size shrinks, naturally leading to a higher error signal that triggers keyframe insertion. Conversely, in static datasets with high overlap, the error remains stable. This mechanism ensures that the system is most responsive when scene coverage is at risk.
@@ 3.2.3 Combined Error Score
1:function ComputeHybridError($I_{t},I_{k},D_{k},\text{Pose}_{t},\text{Pose}_{k},\text{Intrinsics}$)
2:  $\hat{I}_{k},\mathcal{M}\leftarrow\texttt{WarpFrame}(I_{k},D_{k},\text{Pose}_{k},\text{Pose}_{t},\text{Intrinsics})$
@@ 3.4 Hyperparameter Analysis
The core logic of our error computation is summarized in Algorithm 1. In our experiments, these were set as follows based on empirical evaluation on a held-out validation dataset:
  • Error Weights ($\alpha,\beta$): Set to $\alpha=0.7$ and $\beta=0.3$ to prioritize the capture of fine-grained detail. The behavior of our framework is governed by a set of key hyperparameters. The parameters $\alpha$ and $\beta$ in Eq. (3) are treated independently rather than as a single mixing parameter to allow fine-grained control over the system’s sensitivity to sudden lighting changes (captured by SSIM) versus raw pixel intensity shifts.
  • Window Size ($W$): A window size of $W=5$ frames was used. This provides a balance between statistical stability and reactivity to sudden changes in scene dynamics. We employ a small window size $W=5$ for the moving average to prioritize high responsiveness to rapid motion, which a standard Exponential Moving Average (EMA) might over-smooth.
@@
  • Base Threshold ($\theta_{0}$): A small base threshold was set to ensure a minimum level of selectivity even in completely static scenes. The base threshold $\theta_{0}$ was determined through a grid search on the ’Bonn Dynamic’ validation subset, specifically targeting a balance between reconstruction accuracy and frame count.
@@ 4.1 Datasets and Evaluation Metrics
We evaluate reconstruction quality using a suite of standard metrics as in prior works [4, 5, 7, 6]: Accuracy (Acc↓) and Completion (Comp↓) measure the distance between the predicted and ground truth point clouds, Normal Consistency (NC↑) assesses surface smoothness, and Chamfer Distance (CD↓) captures overall shape similarity. To specifically quantify selection efficiency, we introduce the Keyframe Compression Ratio (KFCR↑). This metric represents the percentage of frames discarded from the original sequence, where a higher value indicates greater efficiency and compression. Since our primary contribution is the adaptive selection of the most informative frames for geometry, the KFCR and reconstruction quality serve as the most relevant indicators of performance. We prioritize surface metrics such as Accuracy, Completeness, and Chamfer Distance over pose metrics (ATE/RPE).
@@ 4.2.2 Performance on Dynamic Datasets
On the challenging real-world Bonn dataset, our method achieves superior results across all reconstruction metrics for both Spann3r and CUT3R, while also using fewer keyframes (higher KFCR). This demonstrates the core advantage of our approach: by intelligently selecting frames that capture meaningful changes, it not only improves efficiency but also leads to a more accurate and complete final reconstruction. The ability to filter redundant information allows the reconstruction network to focus its capacity on modeling the most critical aspects of the evolving scene. A reduction in keyframes leads to a direct linear decrease in back-end optimization memory (G-buffer storage) and computational load for the global mapping process, as fewer frames require fusion into the global model.
@@ 4.3 Qualitative Analysis
Figure 2: Qualitative Results on the dynamic BONN, and our own collected outdoor dataset. We compare our method with concurrent works, CUT3R. Our method achieves the best qualitative results in complex and cluttered environments.
Beyond quantitative metrics, we provide a qualitative analysis to inspect the differences in reconstruction quality visually. As illustrated in Figure 2, which shows results on challenging sequences from the BONN dataset, and our own collected outdoor scenes, our method produces visibly more complete and detailed 3D models. When integrated with networks such as CUT3R [6], our adaptive selection helps capture fine details on moving objects and reduces artifacts in the reconstructed geometry. The impact of the module is more significant when evaluated on outdoor scenes, particularly in dynamic environments. This visual evidence supports the quantitative findings and highlights the practical benefits of our approach to real-time robotic perception in dynamic environments.
@@ 4.4.1 Ablation Study
In contrast, our full model integrating both photometric and SSIM errors achieves the best balance of reconstruction quality and efficiency. It outperforms all baselines in accuracy, Chamfer distance, and normal consistency, while maintaining a competitive keyframe compression ratio. This confirms that the complementary integration of photometric sensitivity and structural robustness yields superior keyframe selection performance compared to prior strategies. The ablation study was conducted on the Bonn dataset as it represents the most diverse and challenging dynamic scenarios in our test suite. However, the performance gains observed in Section 4.2 across other datasets confirm that the threshold update dynamics remain robust and effective across various motion profiles.
@@ 4.4.2 Scene-wise Efficiency and Quality
5 Discussion
Assumptions and Limitations: Our approach assumes the availability of depth data (RGB-D) for the warping module. While SSIM provides robustness to global lighting changes, the system may still be sensitive to extreme local specularities or saturated regions. Furthermore, in cases of extremely rapid motion where the overlap mask $\mathcal{M}$ becomes negligible, the warping-based error may lose its geometric grounding, though this typically results in a ’fail-safe’ trigger of a new keyframe.
Modest Gains vs. Efficiency: While the quantitative improvements in Accuracy and Completeness are modest in some sequences, the primary ’win’ of this work is the efficiency gain. We demonstrate that it is possible to achieve parity with or slightly exceed state-of-the-art reconstruction quality while processing significantly fewer frames. This reduction is vital for deploying dense SLAM on resource-constrained robotic platforms.
6 Conclusion and Future Works
@@
The scientific value of this work lies in the formulation of a momentum-aware thresholding strategy that bridges the gap between simple temporal heuristics and complex, computationally expensive keyframe optimization. Practically, this provides a scalable solution for real-time robotic perception, allowing for high-quality environment modeling under restricted bandwidth and memory constraints. For future work, our work on efficient world understanding serves as a foundational step for several exciting research directions. First, while our hybrid error metric is robust, its reliance on appearance makes it sensitive to challenging illumination conditions [25], which we want to address in the future. Second, we plan to leverage the high-fidelity and compact 3D representations created by our method to address the long-tail data problem in robotics. The curated keyframes and resulting 3D models can serve as grounded "digital twins" of real-world events.
{credits}
6.0.1 Acknowledgements
This work was supported by the DARPA YFA Grant D22AP00156-00, the DEVCOM ARL grant SARA W911NF-24-2-0057 grant, and the NSF CPS Grant CNS-2121391.